# SofT-GRPO: Surpassing Discrete-Token LLM Reinforcement Learning via Gumbel-Reparameterized Soft-Thinking Policy Optimization

### v1

@@ 2.3 Attempts of RLVR on Soft-Thinking
In deriving the off-policy soft-thinking probability in GRPO, they restore the value of $\hat{\boldsymbol{s}}_{t}$ in the rollout and calculate the log probability for the soft-thinking reasoning path $\boldsymbol{S}$ with the Gaussian reparameterization trick as follows, keeping the updating process of the answering part $\mathbf{A}$ unchanged:
@@
However, we observe that there are two drawbacks to this method. 1): Although it may help avoid inputs outside the pre-trained discrete-token embedding space, adding noise to inputs instead of logits is not direct and may theoretically mismatch the LLM predictions. As analyzed in Appendix C.1, the embeddings of each token may be linearly dependent, so it becomes hard to attribute which probability $p_{i}$ may contribute to effectiveness. Moreover, the added noise may even be impossible to represent by the embeddings. 2): The methods in (butt2025soft) do not use any advanced sampling methods for effectiveness in training (e.g., incorporating Dirichlet resampling or Gumbel Softmax technique (wu2025llms)), which will undermine the performance of sampled soft-thinking reasoning paths.
So empirically, although observing improvements compared to discrete-token GRPO fine-tuning on the Pass@32 metrics (i.e., the pass rate with 32 attempts), there is a severe degradation on the average accuracy (butt2025soft). As a result, existing attempts of applying GRPO to soft-thinking will not keep its advantage over the discrete-token CoT under the no-finetune setting.
@@
$\Bigg[\sum_{g=1}^{G}\frac{1}{\left|\boldsymbol{S}_{t}\right|+\left|\boldsymbol{A}_{t}\right|}\sum_{t=1}^{\left|\boldsymbol{S}_{t}\right|+\left|\boldsymbol{A}_{t}\right|}\Big(\min\left(p_{g,t}\hat{A}_{g},\text{clip}(p_{g,t},1-\epsilon,1+\epsilon)\hat{A}_{g}\right)-\nabla\beta D_{KL}(\pi_{\theta}||\pi_{\theta_{\text{ref}}})\Bigg]$
@@ 3 SofT-GRPO: Reinforcing Soft-Thinking Policy with Gumbel Reparameterization
To fully unlock the potential of the soft-thinking paradigm and establish a reliable framework to reinforce it, we present an effective RLVR algorithm for soft-thinking policies, SofT-GRPO. As shown in Figure 2, the proposed SofT-GRPO first samples soft-thinking CoTs with controllable randomness using the Gumbel-Softmax technique (wu2025llms; jang2016categorical). In the following policy update stage, we propose a novel SofT-GRPO loss function using the Gumbel reparameterization trick.
@@ 3.1 Group Rollout with Gumbel Noise
As an off-policy RLVR algorithm, for each Query $\boldsymbol{Q}$, SofT-GRPO samples and restores a group of $G$ soft-thinking CoTs in the rollout stage in parallel. As discussed in Section 2.2, advanced sampling methods will bring performance improvement together with stochasticity. So, to explore diverse and powerful soft-thinking reasoning paths, we introduce the Gumbel-Softmax resampling technique (shown in Eq. (4)) in the rollout process as follows:
$(p_{1},\ldots,p_{|\mathcal{T}|})=\pi_{\theta_{\text{old}}}(\cdot|[\boldsymbol{Q},(s_{1},\ldots,s_{t-1})]),$ (9)
@@
where $\boldsymbol{e}_{i}$ is the d-dimensional embedding of the $i$-th token, $\pi_{\theta_{\text{old}}}$ is the old LLM policy used in rollout, and $\epsilon_{i}$ is the Gumbel noise. Using the inverse transform sampling, we can sample $\epsilon_{i}$ by computing $g=-\log(-\log(u))$ where $u\sim\text{Uniform}(0,1)$ (jang2016categorical). Different from adding noise in other distributions, the Gumbel Softmax technique can ensure the stability of the added Gumbel Noise. As shown in Theorem 10 (Refer to Appendix C.2 for proof), the Gumbel Softmax will preserve the distribution of LLM’s output probabilities after adding noise for stochasticity. So, it may simulate the multinomial sampling and reduce the risk of making $\boldsymbol{s}_{t}$ drop outside the pre-trained discrete-token embedding space.
@@
Let $(p_{1},\dots,p_{n})$ be nonnegative, and $\epsilon_{1},\dots,\epsilon_{n}$ independent samples from $\mathrm{Gumbel}(0,1)$ (maddison2016concrete),
$\Pr\left(j=\arg\max_{i}\left(\epsilon_{i}+\log p_{i}\right)\right)=\frac{p_{j}}{\sum_{i=1}^{n}p_{i}}.$ (10)
@@ 3.2 Gumbel Reparameterization for Loss Function
When fine-tuning with the conventional discrete-token GRPO, as represented in Eq. (1), the probability of sampled trajectories can be easily obtained from the output Categorical distributions. However, calculating such a probability is more difficult in soft-thinking (jain2025learning). In the rollout process, we estimate the probability with the reparameterization trick over Gumbel noises as follows:
$\log p(\boldsymbol{s}_{t}|[\boldsymbol{Q},(\boldsymbol{s}_{1},\ldots,\boldsymbol{s}_{t-1})],\theta_{\text{old}})=\sum_{i=1}^{|\mathcal{T}|}-\epsilon_{i}-\exp(-\epsilon_{i}),$ (11)
where $\boldsymbol{s}_{t}$ is built from Eq. (9). When doing the on-policy sampling in SofT-GRPO, we reconstruct the group of soft-thinking reasoning paths $\boldsymbol{S}=(\boldsymbol{s}_{1},\ldots,\boldsymbol{s}_{|\boldsymbol{S}|})$ from the restored $y^{\prime}_{i}$ values in the rollout process in Eq. (9). Then, the log prob is generated with the restored $g^{\prime}_{i}$ values as follows:
$\log p(\boldsymbol{s}_{t}|[\boldsymbol{Q},(\boldsymbol{s}_{1},\ldots,\boldsymbol{s}_{t-1})],\theta)=$ (12)
$\qquad\quad\sum_{i=1}^{|\mathcal{T}|}-(g^{\prime}_{i}-\log p_{i})-\exp(-(g^{\prime}_{i}-\log p_{i})),$
$\text{where}\quad(p_{1},\ldots,p_{|\mathcal{T}|})=\pi_{\theta}(\cdot|[\boldsymbol{Q},(\boldsymbol{s}_{1},\ldots,\boldsymbol{s}_{t-1})]).$
Finally, we can represent the total loss of the proposed SofT-GRPO as shown in Eq. (8)
3.3 Detailed Settings
We follow the general setting in soft-thinking works (wu2025llms), enabling both the top-p and top-k sampling strategies. In both training and inference, we set top-p as $0.95$ and top-k as $5$, the temperature of LLMs $\tau=0.6$, and the temperature in Gumbel Softmax $\tau_{g}=0.1$.
Table 1: Experiment results of baselines and the proposed Latent-RPC on five numerical reasoning benchmarks. We cover 3 base LLMs from 1.5B to 7B and two reasoning patterns, i.e., discrete-token CoT and soft-thinking reasoning. @1 metrics denote the Mean@32 values, where we run each method 32 times on the dataset for average Pass@1 accuracies. @16 and @32 denote the Pass@16 and Pass@32 values on the dataset, respectively. We multiply all the results by 100 to highlight the differences between results. The best result on each metric and dataset is underlined, the best average result is bolded, and the second-best average result is shaded.
@@ 4 Experiments
4.1 Implementation Detail
Training & Testing Settings
We employ DeepScaler (deepscaler2025) as the training dataset, which contains 40,315 queries. In implementing SofT-GRPO, we use SGLang (zheng2024sglang) [^1] Specifically, we modify the rollout process based on the SGLang implementation in https://github.com/eric-ai-lab/Soft-Thinking for the rollout process and the verl-0.4.x framework (sheng2024hybridflow) for policy update.
We involve five famous numerical benchmarks as test sets, i.e., AIME2024, AIME2025, AMC23, MATH-500, and GSM8K. The maximum generation length is confined to 8192 in training & validating, and 32768 in testing. Test answers are verified using the Math Verify package (Kydlicek_Math-Verify_Math_Verification). All experiments in this paper are implemented on a node of 8× NVIDIA H200 GPUs (141 GB VRAM each). With the 64 batch size and a 1e-6 learning rate, running SofT-GRPO on a 1.5B LLM will take approximately 45 hours.
Table 2: Comparison of the proposed SofT-GRPO to the RLVR fine-tuning method proposed in (butt2025soft). Due to the difficulty in reproduction, we compare to their reported results under the same training dataset. GRPO* represents the reported result of GRPO in its paper, and Soft Tokens* represents its reported results fine-tuned under the soft-thinking pattern.
Dataset MATH-500 GSM8K
Metrics @1 @32 @1 @32
LLaMA-3.2-3B-Instruct Base LLM
Discrete-Token CoT Reasoning Pattern
No-Finetune 38.1 84.0 67.9 94.6
+ GRPO 48.3 82.6 79.6 96.5
+ GRPO* 48.3 78.1 79.5 96.6
+ Soft Tokens* 37.4 80.6 71.0 98.0
Soft-Thinking Reasoning Pattern
+ Soft Tokens* 41.3 77.9 75.5 95.2
47.2 83.4 77.6 97.7
+ SofT-GRPO (+5.9) (+5.5) (+2.1) (+2.5)
@@ 4.1 Implementation Detail
Dataset AIME2024 AIME2025 AMC23 GSM8K Average
Metrics @1 M@16 M@32 @1 M@16 M@32 @1 M@16 M@32 @1 M@16 M@32 @1 M@16 M@32
@@
No Fine-tune 30.6 56.7 60.0 23.0 36.7 40.0 70.7 90.0 95.0 81.5 89.5 89.5 51.5 68.2 71.1
+ GRPO 31.8 53.3 50.0 25.3 36.7 30.0 77.3 92.5 92.5 84.9 90.1 90.4 54.8 68.1 65.7
@@
No Fine-tune 27.3 56.7 60.0 23.8 33.3 33.3 69.9 95.0 92.5 81.0 89.3 89.3 50.5 68.6 68.8
+ GRPO 29.2 43.3 46.7 25.4 30.0 30.0 75.8 92.5 92.5 84.9 90.5 90.8 53.8 64.1 65.0
+ SofT-GRPO 32.6 60.0 63.3 26.1 33.3 36.7 76.4 92.5 95.0 85.5 90.6 90.5 55.2 69.1 71.4
Baselines
We include discrete-token GRPO trained on the same dataset and base LLMs without fine-tuning as baselines. In comparing the results of baselines and the proposed SofT-GRPO, we cover two reasoning patterns, i.e., conventional discrete-token CoT reasoning and soft-thinking reasoning. For baselines trained with discrete tokens, we implement their soft-thinking reasoning pattern with the effective Gumbel-Softmax technique (the temperature $\tau_{g}=0.5$) proposed in wu2025llms. Under both patterns, we follow the hyperparameters in (wu2025llms), setting the temperature $\tau$ of LLMs to 0.6 (we discuss this setting in Appendix D.2), top-p to 0.95, and top-k to 30.
Metrics
In evaluating the performances of LLMs, we cover the general metrics of Mean@32, Pass@16, and Pass@32. Mean@32 judges the average Pass@1 accuracy over 32 runs on the datasets. Pass@16 and Pass@32 measure the average probability of covering the accurate answer within 16 and 32 runs, respectively.
@@ 4.2 Main Result
The main result of SofT-GRPO is shown in Table 1. SofT-GRPO can lead to a clear and consistent improvement from the No-Finetune results under the soft-thinking reasoning pattern. Compared to GRPO under the discrete-token CoT, the proposed SofT-GRPO exhibits slight but stable leads (+0.13% on average) on the Mean@32 (i.e., Pass@1, @1 in the Table) metrics over LLMs of three sizes. Notably, the SofT-GRPO can lead to a clear improvement in the Pass@16 (+1.80% on average) and Pass@32 (+2.19%) metrics. Moreover, this observation will still hold when adopting the soft-thinking pattern on the GRPO, demonstrating the significance of SofT-GRPO in reinforcing the soft-thinking performance.
4.3 Comparison to butt2025soft
In this subsection, we conduct comparison experiments with an existing RLVR algorithm for the soft-thinking pattern, the method in butt2025soft. We note this method as Soft Tokens in Table 2, where the proposed SofT-GRPO can demonstrate a clear update compared to the results reported in butt2025soft. Compared to SofT-GRPO, the algorithm in butt2025soft requires transitioning the $d$-dimensional vector between rollout workers and RLVR workers, improving the difficulty of implementation.

### v2

@@ 2.3 Related Work on RLVR Soft-Thinking
In deriving the soft-thinking probability in GRPO, they restore the value of $\hat{\boldsymbol{s}}_{t}$ in the rollout and calculate the log probability for the soft-thinking reasoning path $\boldsymbol{S}$ with the Gaussian reparameterization trick as follows, keeping the updating process of the answering part $\boldsymbol{A}$ unchanged:
@@
Although this method can improve the soft-thinking ability of base LLMs and show improvements compared to discrete-token GRPO on the Pass@32 metrics (i.e., the pass rate with 32 attempts), there is a severe degradation in the average accuracy. This is mainly due to the following two drawbacks: (1) Adding noises on soft-tokens may help avoid inputs outside the pre-trained discrete-token embedding space, but adding noise to soft-tokens instead of logits is not direct and may theoretically mismatch the LLM predictions. As analyzed in Appendix C.1, token embeddings are linearly dependent ($|\mathcal{T}|\gg d$), so they cannot attribute which probability $p_{i}$ is beneficial to the effectiveness of added noise. Moreover, the added noise may even be impossible to be represented by the embeddings. (2) They do not use advanced sampling methods for effectiveness in training (e.g., incorporating Dirichlet resampling or Gumbel-Softmax technique (Wu et al., 2025a)), which will undermine the performance of sampled soft-thinking reasoning paths.
As a result, existing attempts of applying GRPO to soft-thinking will not keep its advantage over the discrete-token CoT under the no-finetune setting.
@@
$\Bigg[\sum_{g=1}^{G}\frac{1}{\left|\boldsymbol{S}_{g}\right|+\left|\boldsymbol{A}_{g}\right|}\sum_{t=1}^{\left|\boldsymbol{S}_{g}\right|+\left|\boldsymbol{A}_{g}\right|}\Big(\min\left(p_{g,t}\hat{A}_{g},\text{clip}(p_{g,t},1-\epsilon,1+\epsilon)\hat{A}_{g}\right)-\beta\mathbb{D}_{KL}(\pi_{\theta}||\pi_{\theta_{\text{ref}}})\Bigg]$
@@ 3 SofT-GRPO: Reinforcing Soft-Thinking Policy with Gumbel Reparameterization
To fully unlock the potential of soft-thinking for better LLM reasoning, we present the RLVR algorithm SofT-GRPO. As shown in Figure 2, it first samples soft-thinking reasoning paths with controllable randomness using the Gumbel-Softmax technique (Wu et al., 2025a). In the following policy update stage, we propose a novel SofT-GRPO loss function with the Gumbel reparameterization trick, which accurately assigns reward credits to token probabilities.
@@ 3.1 Group Rollout with Gumbel Noise
As an off-policy RLVR algorithm, SofT-GRPO samples a group of $G$ soft-thinking CoTs for each query $\boldsymbol{Q}$ during rollout and saves them in a replay buffer. A principal challenge for applying RLVR to soft-thinking is that vanilla soft-thinking generates deterministic soft tokens, as each $\boldsymbol{s}_{t}$ is a fixed weighted sum of embeddings based solely on model probabilities (Eq. (3)). So, to explore diverse and powerful soft-thinking reasoning paths in the rollout process, we introduce the Gumbel-Softmax resampling technique (shown in Eq. (4)) as follows:
$(p_{1},\ldots,p_{|\mathcal{T}|})=\pi_{\theta_{\text{old}}}(\cdot|[\boldsymbol{Q},(\boldsymbol{s}_{1},\ldots,\boldsymbol{s}_{t-1})]),$ (9)
@@
where $\boldsymbol{e}_{i}$ is the embedding of the $i$-th token, $\pi_{\theta_{\text{old}}}$ is the sampling policy used in rollout, and $\epsilon_{i}$ is a scalar Gumbel noise. Using the inverse transform sampling (Jang et al., 2016), we sample $\epsilon_{i}=-\log(-\log(u))$ where $u\sim\text{Uniform}(0,1)$. Besides bringing stochasticity, as discussed in Section 2.2, resampling with Gumbel-Softmax will bring potential improvements on the quality of reasoning paths.
Due to limitations in the amount of fine-tuning data, fine-tuning LLMs with soft-thinking needs to preserve the pre-trained knowledge. So, applying RLVR to soft-thinking should also keep the generated soft tokens $\boldsymbol{s}_{t}$ within the pre-trained discrete-token embedding space. In contrast to noise injection based on other distributions (e.g., Dirichlet or Gaussian), Gumbel-Softmax can better maintain the stability of the soft-token embeddings when providing stochasticity. As established in Theorem 10 (see Appendix D for proof), it produces a relaxation that maintains samples close to the discrete token distribution, thus mitigating the risk of generating out-of-vocabulary soft tokens.
@@
Let $(p_{1},\dots,p_{n})$ be nonnegative, and $\epsilon_{1},\dots,\epsilon_{n}$ independent samples from $\mathrm{Gumbel}(0,1)$ (Maddison et al., 2016),
$\Pr\left(j=\arg\max_{i}\left(\log p_{i}+\epsilon_{i}\right)\right)=\frac{p_{j}}{\sum_{i=1}^{n}p_{i}}.$ (10)
@@ 3.2 Gumbel Reparameterization for Loss Function
In RLVR with discrete tokens (Eq. (1)), the probability of CoT trajectories can be obtained from the output Categorical distributions. Calculating such a probability is impossible in vanilla soft-thinking. To address this, SofT-GRPO proposes using the Gumbel reparameterization trick after obtaining the Gumbel noise $\boldsymbol{\epsilon}$ during the rollout process. We take $\boldsymbol{\epsilon}$ as RLVR actions, consider transforming $\boldsymbol{\epsilon}$ to soft tokens $\boldsymbol{s}_{t}$ as a part of the transition, and estimate the next-soft-token probability with the density of Gumbel noise as follows:
$p(\boldsymbol{g}^{\prime}|[\boldsymbol{Q},(\boldsymbol{s}_{1},\ldots,\boldsymbol{s}_{t-1})],\theta_{\text{old}})=\exp\big[\sum_{i=1}^{|\mathcal{T}|}-\epsilon_{i}-\exp(-\epsilon_{i})\big],$ (11)
where $\boldsymbol{g}^{\prime}=(g^{\prime}_{1},\ldots,g^{\prime}_{|\mathcal{T}|})$ is the $g^{\prime}_{i}$ values in Eq. (9) collected from the replay buffer. We provide the derivation for this equation in Appendix C.3. Following GRPO, SofT-GRPO requires off-policy probabilities for clipping. When calculating off-policy probabilities, we first reconstruct the soft-thinking reasoning paths $\boldsymbol{S}=(\boldsymbol{s}_{1},\ldots,\boldsymbol{s}_{|\boldsymbol{S}|})$ from the restored $y^{\prime}_{i}$ values in the rollout process. Then, the probability is calculated with the restored $g^{\prime}_{i}$ values as follows:
$p(\boldsymbol{g}^{\prime}|[\boldsymbol{Q},(\boldsymbol{s}_{1},\ldots,\boldsymbol{s}_{t-1})],\theta)=$ (12)
$\qquad\exp\left(\sum_{i=1}^{|\mathcal{T}|}-(g^{\prime}_{i}-\log p_{i})-\exp(-(g^{\prime}_{i}-\log p_{i}))\right),$
$\text{where}\ (p_{1},\ldots,p_{|\mathcal{T}|})=\pi_{\theta}(\cdot|[\boldsymbol{Q},(\boldsymbol{s}_{1},\ldots,\boldsymbol{s}_{t-1})]).$
Finally, the SofT-GRPO loss function is Eq. (8), where $\pi$ and $\pi_{\text{ref}}$ are token probability distributions that take history soft-tokens $\boldsymbol{S}$ and answer tokens $\boldsymbol{A}$ as input. $p_{g,t}$ is calculated as the probability ratio between current policy $p_{\theta}$ (Eq. (11)) and sampling policy $p_{\theta_{\text{old}}}$ (Eq. (12)). Intuitively, SofT-GRPO recalculates the noise corresponding to $\boldsymbol{g}^{\prime}$ obtained in rollout and encourages this noise only when the current soft-thinking trajectory leads to a higher $\hat{A}_{g}$. Unlike discrete-token GRPO in Eq. (2), SofT-GRPO updates the probability of several tokens at each LLM step $t$.
Table 1: Experiment results of baselines and the proposed SofT-GRPO on five numerical reasoning benchmarks. We cover 3 LLMs from 1.5B to 7B and two reasoning patterns, i.e., discrete-token CoT and soft-thinking reasoning. @1 metrics denote the Mean@32 values, where we run each method 32 times on the dataset for average Pass@1 accuracies. @16 and @32 denote the Pass@16 and Pass@32 values on the dataset, respectively. We multiply all the results by 100 to highlight the differences between results. The best result on each metric and dataset is underlined, the best average result is bolded, and the second-best average result is shaded.
@@ 4 Experiments
4.1 Implementation Details
Detailed Settings. In practice, we follow the general setting in soft-thinking works (Wu et al., 2025a), enabling both the top-p and top-k sampling strategies, considering only tokens with $k$ highest probabilities instead of the whole token set $|\mathcal{T}|$ and normalizing their probabilities. In both training and inference, we set top-p as $0.95$ and top-k as $5$, the temperature of LLMs $\tau=1$, and the temperature in Gumbel-Softmax $\tau_{g}=0.1$.
Training & Testing Settings. We employ DeepScaler (Luo et al., 2025) as the training dataset, which contains 40,315 queries. In implementing SofT-GRPO, we use SGLang (Zheng et al., 2024) for the rollout process and the verl-0.4.x framework (Sheng et al., 2024) for RLVR.
We involve five numerical reasoning benchmarks as in-domain test sets (AIME2024, AIME2025, AMC23, MATH-500, and GSM8K). We also employ a scientific reasoning benchmark (GPQA Diamond) and two code benchmarks (HumanEval and MBPP) for out-of-domain evaluation. The maximum generation length is 8192 in training and 32768 in testing. Test answers are verified using the Math Verify package (Kydlíček, 2025). All experiments are implemented on a node of 8× NVIDIA H200 GPUs (141 GB VRAM each), and detailed parameters are shown in Appendix D.
@@ 4.1 Implementation Details
Dataset AIME2024 AIME2025 AMC23 MATH-500 GSM8K Average
Metrics @1 M@16 M@32 @1 M@16 M@32 @1 M@16 M@32 @1 M@16 M@32 @1 M@16 M@32 @1 M@16 M@32
@@
No Fine-tune 30.6 56.7 60.0 23.0 36.7 40.0 70.7 90.0 95.0 84.6 92.6 92.4 81.5 89.5 89.5 58.1 73.1 75.4
+ GRPO 31.8 53.3 50.0 25.3 36.7 30.0 77.3 92.5 92.5 87.1 91.6 91.4 84.9 90.1 90.4 61.3 72.8 70.9
@@
No Fine-tune 27.3 56.7 60.0 23.8 33.3 33.3 69.9 95.0 92.5 79.4 88.8 89.8 81.0 89.3 89.3 56.3 72.6 73.0
+ GRPO 29.2 43.3 46.7 25.4 30.0 30.0 75.8 92.5 92.5 86.3 91.0 91.4 84.9 90.5 90.8 60.3 69.5 70.3
+ SofT-GRPO 32.6 60.0 63.3 26.1 33.3 36.7 76.4 92.5 95.0 86.3 92.2 92.0 85.5 90.6 90.5 61.4 73.7 75.5
Baselines. We use both conventional discrete-token CoT reasoning and soft-thinking reasoning as baselines for SofT-GRPO. For discrete-token CoT, we include discrete-token GRPO trained on the same dataset and base LLMs without fine-tuning as baselines. For baselines with soft-thinking, we implement the soft-thinking reasoning pattern with the Gumbel-Softmax technique, using the best temperature setting $\tau_{g}=0.5$ proposed in Wu et al. (2025a). Under both patterns, we follow the parameters in (Wu et al., 2025a), setting the temperature $\tau$ of LLMs to 0.6 (we discuss this setting in Appendix E.2), top-p to 0.95, and top-k to 30.
Metrics. We adopt the general metrics of Mean@32 (the average Pass@1 accuracy over 32 runs), Pass@16, and Pass@32, which measure the average probability of covering the correct answer within 16 and 32 runs, respectively.
@@ 4.2 Main Result
As shown in Table 1, SofT-GRPO can lead to a clear and consistent improvement from the No-Finetune results under the soft-thinking reasoning pattern. Moreover, SofT-GRPO outperforms applying soft-thinking on top of the discrete-token trained GRPO on average, demonstrating superior soft-thinking reasoning ability.
For Pass@1, SofT-GRPO consistently outperforms the discrete-token trained GRPO across LLMs of three sizes (+0.13% on average accuracy). In contrast, comparing the +GRPO results on the discrete-token and soft-thinking patterns, performing soft-thinking on top of the discrete-token trained GRPO gives inconsistent improvement.
As observed in Yue et al. (2025), GRPO can lead to decreased performance on Pass@K over the No-Finetune model, even though it improves performance for Pass@1. We can observe the same result in Table 1. In contrast, the SofT-GRPO leads to a clear improvement in the Pass@16 (+1.80% on average) and Pass@32 (+2.19% on average) metrics. Appendix E.2 shows that the superiority will also hold for different temperatures. As detailed in Appendix F, the improvement in Pass@16 and Pass@32 is likely due to the fact that for each sample in each LLM reasoning step, SofT-GRPO can reinforce the output probability of several tokens instead of focusing on one in GRPO.
4.3 Comparison in Out-of-Domain Datasets
Besides numerical reasoning, we conduct out-of-domain experiments to evaluate the general reasoning ability of SofT-GRPO. As shown in Table 2, the LLMs fine-tuned with SofT-GRPO on numerical questions can still demonstrate advantages from Pass@1 to Pass@32 on a scientific reasoning benchmark (GPQA Diamond) and two code benchmarks (HumanEval and MBPP).
4.4 Boosting SofT-GRPO with Majority Voting
To further exploit the advantage of SofT-GRPO on Pass@32, in this subsection, we design to boost SofT-GRPO with majority voting (Chen et al., 2024). As shown in Table 3, SofT-GRPO with majority-voting can outperform No-Finetune LLM and LLM fine-tuned with discrete-token GRPO across Major@16 (the accuracy of the most common answer in 16 runs) and Major@32. The results show that SofT-GRPO-fine-tuned LLMs can be strengthened into better reasoning solvers with majority voting.
Table 4: Comparison of the proposed SofT-GRPO to the RLVR fine-tuning method proposed in (Butt et al., 2025). We compare to their reported results under the same training dataset and LLM. Soft Tokens* represents their reported results fine-tuned under the soft-thinking pattern.
Dataset MATH-500 GSM8K
Metrics @1 @32 @1 @32
LLaMA-3.2-3B-Instruct Base LLM
Discrete-Token CoT Reasoning Pattern
No-Finetune 38.1 84.0 67.9 94.6
+ GRPO 48.3 82.6 79.6 96.5
Soft-Thinking Reasoning Pattern
+ Soft Tokens* 41.3 77.9 75.5 95.2
47.2 83.4 77.6 97.7
+ SofT-GRPO (+5.9) (+5.5) (+2.1) (+2.5)
4.5 Comparison to Butt et al. (2025)
We conduct comparison experiments with an existing RLVR algorithm for the soft-thinking pattern, the method in Butt et al. (2025). We refer to this method as Soft Tokens in Table 4, where the proposed SofT-GRPO can demonstrate a clear improvement compared to the results reported in Butt et al. (2025). Compared to SofT-GRPO, the algorithm in Butt et al. (2025) also requires transitioning the $d$-dimensional vector between rollout workers and RLVR workers, making implementation more difficult.